OpenStudy (anonymous):

Peter drove 300 miles in 5 hours. Jane drove 400 miles in 8 hours. Who drove at a higher speed? A. Peter B. Jane

OpenStudy (anonymous):

\[V=\frac{ S }{ T }\]Peter speed:\[\frac{ 300 }{ 5 } = 60 m/h\]Jane speed:\[\frac{ 400 }{ 8 } = 50m/h\]So, Peter drove faster.

OpenStudy (anonymous):

miles per hour means \(\frac{\text{miles }}{\text{hours}}\)